# Structural characterizations and Obstructions to Planar-Rips complexes

### v1

@@ 3. Classification of Rips complexes with weak-pseudomanifold structure
In this section, we present a complete classification of well-behaved spaces known as pseudomanifolds and weak-pseudomanifolds that admit planar-Rips structure. To accomplish this, we first introduce the necessary terminologies and tools before proceeding with the proof of A by induction. Subsequently, we use this to further prove B.
The dimension of a simplex will be one less than its cardinality. The dimension of a simplicial complex is defined to be equal to the supremum of all $n$’s such that $K$ has a face of dimension $n$. A simplex in $K$ is called maximal or a facet if it is not contained in another larger simplex. An $n$-dimensional simplicial complex is said to be pure if every maximal simplex is of dimension $n$.
@@
A pure simplicial complex is said to be strongly connected if for any two facets $\sigma$ and $\gamma$, there exists a sequence of $n$-simplices, say $\lambda_{1},\lambda_{2},\dots\lambda_{k},$ such that $\lambda_{1}=\sigma,\ \lambda_{k}=\gamma$ and $\lambda_{i}\cap\lambda_{i+1}$ is an $(n-1)$-simplex for all $11\leq i\leq k-$ 1.
@@
  • (a)
    An $n$-dimensional simplicial complex is said to be a weak pseudomanifold if it is pure and each $(n-1)$-simplex is contained in exactly two $n$-simplices.
  • (b)
@@
  • (c)
    An $n$-dimensional pseudomanifold is said to be a normal pseudomanifold if the link of each simplex of codimension atleast two is connected.
@@
[2, Theorem 6.3] : For any $d\geq 2$ and any finite subset $X\subset\mathbb{R}^{2}$, if the planar-Rips complex $\mathcal{R}(X)$ is a $d$-dimensional normal pseudomanifold then $\mathcal{R}(X)$ is isomorphic to boundary complex of the $(d+1)$ dimensional cross-polytope.
@@
For $X\subset\mathbb{R}^{2}$, suppose the planar-Rips complex $\mathcal{R}(X)$ is an $n$-dim weak-pseudomanifold where $n\geq 2$, then the link of any simplex of codimension at least two can have at most two strongly connected components.
@@
On contrary, suppose there are more than two strongly connected components in the link of some $k$-simplex, say $\sigma$, where $20\leq k\leq n-$ 2 . We have that the link of the simplex $\sigma$ is again a weak-pseudomanifold of dimension $n-k-1$. Since each strong component is a pseudomanifold and admits a planar-Rips structure, choose two non-adjacent vertices from each of them, hence forming an induced subgraph $K_{1,6}$ This is not possible owing to Lemma 2.5. ∎
@@
For $X\subset\mathbb{R}^{2}$, if $\mathcal{R}(X)$ is an $n$-dim weak-pseudomanifold ($n\geq 2$), then the link of any $(n-2)$ simplex can have at most 11 vertices. Further, it is either a cycle or a disjoint union of exactly two cycles.
@@
Lemma 3.6 (Classification of $2$–dimensional weak-pseudomanifolds).
Any compact weak-pseudomanifold of dimension $2$ consists of a collection of compact $2$–manifolds without boundary glued or self–glued along a finite number of points. More technically: such a weak-pseudomanifold $X$ is the quotient space of a disjoint union of compact 2–manifolds without boundary such that if $q$ is the quotient map, then all points $x\in X$ satisfy #$\#\left\{q^{-1}(x)\right\}<\infty$ and for all but finitely many $x\in X$, #$\#\left\{q^{-1}(x)\right\}=1$.
@@
(a) $\mathcal{R}(X)$: Pinched Torus.
(b) $\mathcal{R}(X-v)$: After deleting the singular vertex.
Figure 2. A sphere with self-identification of two points is homeomorphic to pinched torus. The link of its singular vertex $v$ has two connected components. Here, both $\mathcal{R}(X)$ and $\mathcal{R}(X-v)$ do not admit planar-Rips structure.
By Theorem 2.3, the fundamental group of $\mathcal{R}(X)$ is free. Hence, by the structure theroem for (weak) pseudomanifolds (Lemma 3.6), the pseudomanifold $\mathcal{R}(X)$ is homeomorphic to the quotient space of sphere $\mathbb{S}^{2}$ by taking a self-identification of finite sets of points; let $q:\mathbb{S}^{2}\to\mathcal{R}(X)$ be the associated quotient map.
Let $\tilde{X}$ be the set of singular vertices (vertices whose links are not connected) in $\mathcal{R}(X)$. For any vertex $v\in X$, since $Lk(v)$ has at most two disjoint cycles (by Proposition 3.4), the set $q^{-1}(v)$ can have at most two points, and $q^{-1}(u)\cap q^{-1}(v)=\emptyset$ for $u\neq v$; hence, $|q^{-1}(\tilde{X})|=2|\tilde{X}|$. Now, consider the planar-Rips complex $\mathcal{R}(X-\tilde{X})$ which is homotopy equivalent to a sphere with $2m$ points deleted, where $|\tilde{X}|=m$. Each removed point $v_{i}$ from the sphere corresponds to a cycle $\mathcal{C}_{i}$ in the link of the singular vertex $q(v_{i})$ in $\mathcal{R}(X)$. Figure 2 shows an illustration for $m=1$ case. Further, the cycle $[p(\mathcal{C}_{i})]$ is non-trivial in $\pi_{1}(\mathcal{S}(X-\tilde{X}))$ since the projection map $p:\mathcal{R}(X-\tilde{X})\to\mathcal{S}(X-\tilde{X})$ induces a $\pi_{1}$-isomorphism (Theorem 2.3).
From Proposition 2.8, the convex hulls of any two of the $2m$ cycles in $\mathcal{R}(X-\tilde{X})$ cannot intersect; hence the shadow $\mathcal{S}(X-\tilde{X})$ is homotopy equivalent to wedge of $2m$ copies of $\mathbb{S}^{1}$. This is possible only if $m=0$ since for $m\geq 1$, the Rips complex $\mathcal{R}(X-\tilde{X})$ is homotopy equivalent to a wedge of $(2m-1)$ copies of $\mathbb{S}^{1}$. Therefore, the link of every vertex is connected and the conclusion follows by the virtue of Theorem 3.3.
@@
: Given the boundary complex of $(n+1)$-dimensional cross-polytope, say $\mathcal{O}_{n}$, the antipodal point of any vertex $v\in\mathcal{O}_{n}$ is the vertex in $\mathcal{O}_{n}$ not adjacent to $v$ and is denoted by $\bar{v}$.
@@
For $X\subset\mathbb{R}^{2}$, let $|X|=2n$ and $\mathcal{R}(X)$ be isomorphic to the boundary complex of $n$-dimensional cross-polytope($n\geq 2$). Then any $x\in X$ is not in the convex hull of the remaining vertices i.e. the shadow $\mathcal{S}(X)$ is a simple convex polygon with $2n$ sides.
@@
Let if possible, there exists $x\in X$ such that $x\in conv(X-x)$. Then by Caratheodory’s theorem, there exists vertices $a,b,c\in(X-x)$ such that $x\in\mathrm{conv}\left\{a,b,c\right\}$. Now there are two cases to consider.
  1. (1)
    Suppose one of the points $a,b,c$ is an antipodal point of $x$; without loss of generality, assume $a$ to be the antipodal point $\bar{x}$. Since $d(\bar{x},b)<1$ and $d(\bar{x},c)<1$ in $\mathcal{R}(X)$, the distance between $\bar{x}$ and any point in $\triangle\bar{x}bc$ is less than 1 i.e. $d(\bar{x},x)<1$ which is absurd.
  2. (2)
    Suppose none of the points $a,b,c$ is the antipodal point $\bar{x}$. Then from Lemma 2.6, we have $x\in Lk(\tilde{x})$ i.e. $d(\bar{x},x)<1$ which is again absurd.
@@
[2, Lemma 6.1] Let $\mathcal{S}$ be a simple polygon in $\mathbb{R}^{2}$ and $v$ be a point outside or on $\mathcal{S}$. Suppose further that for every vertex $P$ of $\mathcal{S}$, all the intersections of the ray $vP^{\rightarrow}$ with $\mathcal{S}$ lie between $v$ and $P$. Then there is an edge $AB$ of $\mathcal{S}$ such that for any vertex $P$ of $\mathcal{S}$ the segments $vP$ and $AB$ have non-empty intersection.
We refer to such a point $v$ as guard point. The motivation to use this terminology stems from the Art Gallery problem, a prominent concept in Computational Geometry. In the art gallery problem [9, p. 8], a guard point (or sometimes called point guard) refers to a point on or inside a polygon from which the entire interior of the polygon can be observed. It is a location from where a guard can have an unobstructed line of sight to every point inside the gallery. In our case, the guard points are considered to be lying strictly outside the polygon, and the guard points are needed to have every line of sight passing through an edge of the polygon. Furthermore, from Lemma 3.10, it is sufficient for the guard point to have a clear line of sight to only the vertices of the simple polygon.
@@
3.1. Position of a point relative to a line:
@@
The union of line $L_{AB}$ and the region $L^{+}_{AB}$ will be denoted by $cl(L^{+}_{AB})$, which is same as taking the closure of $L^{+}_{AB}$ in usual topological sense. Note that the order of the points $A$ and $B$ is significant while determining the position of point $P$ relative to the line $L_{AB}$. The orientation of the line is taken to be in the direction from $A$ to $B$. Moreover, it can be easily checked that $L^{+}_{AB}=L^{-}_{BA}$.
@@
=$\mathcal{H}=\left\{X\in\mathbb{R}^{2}\mid C\in\mathrm{conv}\left\{A,B,X\right\}\right\}$
@@
    $\mathcal{H}=cl(L^{+}_{AC})\cap cl(L^{-}_{BC})$, when $A,B,C$ are in counterclockwise order
Figure 3. The two cases of Lemma 3.12: The gray region(extended) denotes the set $cl(L^{-}_{AC})\cap cl(L^{+}_{BC})$ or $cl(L^{+}_{AC})\cap cl(L^{-}_{BC})$, depending on whether points are in clockwise or counterclockwise order respectively.
@@ 3.2. Perpendicular bisector
Given any two points $a$ and $b$ in $\mathbb{R}^{2}$, the line of perpendicular bisector of $a$ and $b$ will be denoted by $L_{a|b}$ and the set of all points whose distance from $a$ is strictly less than distance from $b$ will be denoted by $\mathcal{P}_{a<b}$ i.e.
$\mathcal{P}_{a<b}:=\left\{x\in\mathbb{R}^{2}\mid d(a,x)<d(b,x)\right\}$
Note that the set $\mathcal{P}_{a<b}$ inside $\mathbb{R}^{2}$ is same as the set of all points that lie on the same side of the line $L_{a|b}$ as the point $a$ does i.e. $\mathcal{P}_{a<b}=L^{+}_{a|b}$. Moreover, it can be easily checked that $L^{+}_{A|B}=L^{-}_{B|A}$.
@@
Let $A,B,C,D$ be four points in $\mathbb{R}^{2}$ such that no three points are collinear, the segments $\overline{AB}$ and $\overline{CD}$ intersect, and the distance of $D$ from $C$ is less than the distance from both $A$ and $B$. Suppose $P$ is any point whose distance from $C$ is greater than the distance from both $A$ and $B$, then the point $D$ is in the convex hull of points $A,B,\ \text{and}\ P$ i.e. $D\in\mathrm{conv}\left\{A,B,P\right\}$.
@@
Figure 4. The blue and green regions denote the respective possible locations for points $D$ and $P$ corresponding to a specific arrangement of points $A$, $B$ $C$ satisfying Lemma 3.13. The point $P$ lies in the region of points that satisfy $D\in\mathrm{conv}\left\{A,B,P\right\}$.
Since the segments $\overline{AB}$ and $\overline{CD}$ intersect, and no three points are collinear, the points $C$ and $D$ lie on either side of the line $L_{AB}$; assume without loss of generality, $C\in L^{-}_{AB}$ and $D\in L^{+}_{AB}$. Further, the point $D$ being closer to $C$ than both $A$ and $B$ implies $D\in L^{-}_{A|C}\cap L^{+}_{C|B}$. Likewise, $P\in L^{+}_{A|C}\cap L^{-}_{C|B}$. Now it can be easily checked that $L^{+}_{A|C}\cap L^{-}_{C|B}\subset L^{+}_{AD}\cap L^{-}_{BD}$ i.e. $P\in L^{+}_{AD}\cap L^{-}_{BD}$ (See Figure 4 for an illustration). Consequently, $D\in\mathrm{conv}\left\{A,B,P\right\}$ as follows from Lemma 3.12.
@@
Let $r>0$. Suppose $A,B,C,D,X,Y$ are six points in $\mathbb{R}^{2}$ satisfying the following:
@@
    $ABCD$ is a quadrilateral with sides of length less than $r$ and diagonals of length at least $r$
@@
    The point $X$ is a guard point of $ABCD$ along the side $AB$, and distance of $X$ from the points $C$ and $D$ is less than $r$.
@@
    The distance of $Y$ from points $A$ and $B$ is less than $r$, and distance of $Y$ from $X$ is at least $r$.
Then the point $Y$ is a guard point of $ABCD$ along the side $CD$. Further, the distance of $Y$ from points $C$ and $D$ is less than $r$.
@@
Figure 5. An illustrative example for Lemma 3.14: A possible arrangement of points $A,\ B,\ C,\ D$ and $X$ satisfying the hypotheses. Corresponding to this arrangement, the point $Y$ is located in the gray region.
From $(1)$, it follows that the quadrilateral $ABCD$ is simple and convex. The points $A$, $B$, and $X$ are not collinear as follows from $(2)$ and $(3)$. Hence, without loss of generality, assume the points $C$ and $D$ are on the left side and $X$ on the right side of the line $L_{AB}$. The point $Y$ being closer to both $A$ and $B$ than $X$ implies $Y\in L^{+}_{A|X}\cap L^{-}_{X|B}$ (An illustration of a possible arrangement is shown in Figure 5). We have the following sequential observations:
  1. (a).
    $C\in L^{-}_{A|X}$ since the point $C$ is closer to $X$ than $A$. Likewise, $D\in L^{+}_{X|B}$ since the point $D$ is closer to $X$ than $B$.
  2. (b).
@@
  3. (c).
    As a consequence, $Y\in L^{-}_{CD}$ since $Y\notin\mathrm{conv}\{A,B,C,D,X\}$, given that $d(X,Y)\geq r$.
Hence the point $Y$ lies in the exterior region of side $CD$. From hypothesis $(3)$, we have $d(Y,B)<d(D,B)$ implying that $Y$ lies on right side of the line $L_{BD}$ i.e. $Y\in L^{-}_{BD}$. Likewise, $Y\in L^{+}_{AC}$. Further, the points $C$ and $D$ cannot be inside $\triangle YAB$ since $d(B,D)\geq r$. This adds further restriction on the location where the point $Y$ can lie as apparent in Figure 5. Therefore, the point $Y$ is guard point of $ABCD$ along the side $CD$. Consequently, it follows from Lemma 2.4 that the distance of $Y$ from both $C$ and $D$ is less than $r$.
The above lemma has a further implication that $X$ is a guard point along the side $AB$ if and only if $Y$ is a guard point along the side $CD$. Note that the planar-Rips complex induced by the six points in the above lemma is nothing but the boundary complex of octahedron $\mathcal{O}_{2}$.
@@
We proceed by induction. Firstly, for the case when $n=2$, it follows from Theorem 3.7 that $\mathcal{R}(X)$ is isomorphic to $\mathcal{O}_{2}$. Now let $n>2$ and consider a strongly connected component of link of any $x\in X$, say $\widetilde{\mathcal{O}}$. By induction, the subcomplex $\widetilde{\mathcal{O}}$ is isomorphic to $\mathcal{O}_{n-1}$. From Lemma 3.9, the shadow $\mathcal{S}(\widetilde{\mathcal{O}})$ is a $2n$-sided convex polygon.
@@
The point $x$ cannot be inside the polygon $\mathcal{S}(\widetilde{\mathcal{O}})$
@@
On contrary, suppose $x$ is inside $\mathcal{S}(\widetilde{\mathcal{O}})$. Then there exists an $(n-1)$-dimensional simplex $\Delta$ in $\widetilde{\mathcal{O}}$ such that $x\in\mathrm{conv}((\Delta))$ Since $\mathcal{R}(X)$ is an $n$-dimensional pseudomanifold, the $(n-1)$-dimensional simplex $\Delta$ is shared by exactly two $n$-dimensional simplices, one of which is $\Delta\cup\left\{x\right\}$. Hence, there exists another point $y$ outside $\widetilde{\mathcal{O}}$ such that the vertices of $\Delta$ are adjacent to $y$ i.e. $\Delta\cup\left\{y\right\}$ is an $n$-dimensional simplex in $\mathcal{R}(X)$. Now from Lemma 2.6, we have $x\in Lk(y)$ i.e. $\Delta\cup\left\{x\right\}\cup\left\{y\right\}$ is an $(n+1)$-dimensional simplex. But this is not possible since $\mathcal{R}(X)$ is $n$-dimensional.
Hence the point $x$ is outside the polygon $\mathcal{S}(\widetilde{\mathcal{O}})$. Now we show that $x$ is a guard point of polygon $\mathcal{S}(\widetilde{\mathcal{O}})$ along a side.
@@
There exists a side $ab$ of the polygon $\mathcal{S}(\widetilde{\mathcal{O}})$ such that $x$ is a guard point of $\mathcal{S}(\widetilde{\mathcal{O}})$ along the side $ab$

### v2

@@ 3. Classification of Rips complexes with weak-pseudomanifold structure
In this section, we present a complete classification of well-behaved spaces known as pseudomanifolds and weak-pseudomanifolds admitting a planar-Rips structure. To accomplish this, we first introduce the necessary terminologies and tools before proceeding with the proof of Theorem A by induction. Subsequently, we use this to further prove Theorem B.
The dimension of a simplex will be one less than its cardinality. The dimension of a simplicial complex is defined to be equal to the supremum of all $n$’s such that $K$ has a face of dimension $n$. A simplex in $K$ is called maximal or a facet if it is not contained in another larger simplex. An $n$-dimensional simplicial complex is said to be pure if every maximal simplex is of dimension $n$. We refer the reader to look at [9, 23] for more on these topics.
@@
A pure simplicial complex is said to be strongly connected if for any two facets $\sigma$ and $\gamma$, there exists a sequence of $n$-simplices, say $\lambda_{1},\lambda_{2},\dots\lambda_{k},$ such that $\lambda_{1}=\sigma,\ \lambda_{k}=\gamma$ and $\lambda_{i}\cap\lambda_{i+1}$ is an (n1)(n-1)-simplex for all $1\leq i\leq k-1$.
@@
  1. (a)
    An $n$-dimensional simplicial complex is said to be a weak pseudomanifold if it is pure and each (n1)(n-1)-simplex is contained in exactly two $n$-simplices.
  2. (b)
@@
  3. (c)
    An $n$-dimensional pseudomanifold $(n\geq 2)$ is said to be a normal pseudomanifold if the link of each simplex of codimension at least two is connected.
@@
[3, Theorem 6.3] For any $n\geq 2$ and any finite subset $X\subset\mathbb{R}^{2}$, if the planar-Rips complex $\mathcal{R}(X)$ is an $n$-dimensional normal pseudomanifold then $\mathcal{R}(X)$ is isomorphic to boundary complex of the $(n+1)$ dimensional cross-polytope.
@@
For $X\subset\mathbb{R}^{2}$, suppose the planar-Rips complex $\mathcal{R}(X)$ is an $n$-dimensional weak-pseudomanifold where $n\geq 2$, then the link of any simplex of codimension at least two can have at most two strongly connected components.
@@
On contrary, suppose there are more than two strongly connected components in the link of some $k$-simplex, say $\sigma$, where $0\leq k\leq n-2$ . We have that the link of the simplex $\sigma$ is again a weak-pseudomanifold of dimension nk1n-k-1. Since each strong component is a pseudomanifold and admits a planar-Rips structure, choose two non-adjacent vertices from each of them, hence forming an induced subgraph $K_{1,6}$ . This is not possible owing to Lemma 2.5. ∎
@@
For $X\subset\mathbb{R}^{2}$, if $\mathcal{R}(X)$ is an $n$-dimensional weak-pseudomanifold ($n\geq 2$), then the link of any (n2)(n-2) simplex can have at most 11 vertices. Further, it is either a cycle or a disjoint union of exactly two cycles.
@@
Lemma 3.6 (Classification of two–dimensional weak-pseudomanifolds).
Any compact two-dimensional weak-pseudomanifold $X$ is homeomorphic to the quotient space of a disjoint union of compact 2–manifolds without boundary such that if $q$ is the quotient map, then $\left\{q^{-1}(x)\right\}<\infty$ for all points $x\in X$, and $\left\{q^{-1}(x)\right\}=1$ for all but finitely many $x\in X$.
@@
Figure 2. (On the left) A sphere with self-identification of two points($m=1$) is homeomorphic to pinched torus. The link of its singular vertex $\mathrm{v}$ has two connected components. (On the right) We have induced subcomplex of pinched torus after deleting the singular vertex. Both the structures do not admit planar-Rips structure.
By Theorem 2.3, the fundamental group of $\mathcal{R}(X)$ is free. Hence, by the structure theroem for (weak) pseudomanifolds (Lemma 3.6), the pseudomanifold $\mathcal{R}(X)$ is homeomorphic to the quotient space of sphere $\mathbb{S}^{2}$ by taking a self-identification of finite sets of points; let $q\colon\mathbb{S}^{2}\to\mathcal{R}(X)$ be the associated quotient map.
[…]
@@
Given the boundary complex of $(n+1)$-dimensional cross-polytope, say $\mathcal{O}_{n}$, the antipodal point of any vertex $v\in\mathcal{O}_{n}$ is the vertex in $\mathcal{O}_{n}$ not adjacent to $v$ and is denoted by $\bar{v}$.
@@
For $X\subset\mathbb{R}^{2}$ with $(2n+2)$ points($n\geq 1)$, suppose the planar-Rips complex $\mathcal{R}(X)$ is isomorphic to $\mathcal{O}_{n}$. Then any $x\in X$ is not in the convex hull of the remaining vertices i.e., the shadow $\mathcal{S}(X)$ is a simple convex polygon with $2n$ sides.
@@
Let if possible, there exists $x\in X$ such that $x\in conv(X-x)$. Then it follows by Caratheodory’s theorem that there exists vertices $a,b,c\in(X-x)$ such that $x\in\mathrm{conv}\left\{a,b,c\right\}$. Now there are two cases to consider.
  1. Case 1.
    Suppose one of the points $a,b,c$ is an antipodal point of $x$; without loss of generality, assume $a$ to be the antipodal point $\bar{x}$. Since $d(\bar{x},b)<1$ and $d(\bar{x},c)<1$ in $\mathcal{R}(X)$, the distance between $\bar{x}$ and any point in $\triangle\bar{x}bc$ is less than 1 i.e., $d(\bar{x},x)<1$ which is absurd.
  2. Case 2.
[…]
@@
[3, Lemma 6.1] Let $\mathcal{S}$ be a simple polygon in $\mathbb{R}^{2}$ and $v$ be a point outside or on $\mathcal{S}$. Suppose further that for every vertex $P$ of $\mathcal{S}$, all the intersections of the ray $\overrightarrow{vP}$ with $\mathcal{S}$ lie between $v$ and $P$. Then there is an edge $AB$ of $\mathcal{S}$ such that for any vertex $P$ of $\mathcal{S}$ the segments $vP$ and $AB$ have non-empty intersection.
We refer to such a point $v$ as guard point. The motivation to use this terminology stems from the Art Gallery problem (See [22]), a prominent concept in Computational Geometry. In the art gallery problem [15, p. 8], a guard point (or sometimes called point guard) refers to a point on or inside a polygon from which the entire interior of the polygon can be observed. It is a location from where a guard can have an unobstructed line of sight to every point inside the gallery. In our case, the guard points are considered to be lying strictly outside the polygon, and the guard points are needed to have every line of sight passing through an edge of the polygon. Furthermore, from Lemma 3.10, it is sufficient for the guard point to have a clear line of sight to only the vertices of the simple polygon.
@@
3.1. Position of a point relative to a line.
@@
[…]
@@
$\mathcal{H}=\left\{X\in\mathbb{R}^{2}\mid C\in\mathrm{conv}\left\{A,B,X\right\}\right\}$
@@
    $\mathcal{H}=\mathrm{cl}(L^{+}_{AC})\cap\mathrm{cl}(L^{-}_{BC})$, when $A,B,C$ are in counterclockwise order
[…]
@@ 3.2. Perpendicular bisector
Given any two points $A$ and $B$ in $\mathbb{R}^{2}$, the line of perpendicular bisector of the segment $\overline{AB}$ will be denoted by $L_{A|B}$ and the set of all points whose distance from $A$ is strictly less than distance from $B$ will be denoted by $\mathcal{P}_{A<B}$ i.e.,
$\mathcal{P}_{A<B}\coloneq\left\{x\in\mathbb{R}^{2}\mid d(A,x)<d(B,x)\right\}$
Note that the set $\mathcal{P}_{A<B}$ inside $\mathbb{R}^{2}$ is same as the set of all points that lie on the same side of the line $L_{A|B}$ as the point $A$ does i.e., $\mathcal{P}_{A<B}=L^{+}_{A|B}$. Moreover, it can be easily checked that $L^{+}_{A|B}=L^{-}_{B|A}$.
@@
Let $A,B,C,D$ be four points in $\mathbb{R}^{2}$ such that no three points are collinear, the segments $\overline{AB}$ and $\overline{CD}$ intersect, and the distance of $D$ from $C$ is less than the distance from both $A$ and $B$. Suppose $P$ is another point whose distance from $C$ is greater than the distance from both $A$ and $B$, then the point $D$ is in the convex hull of points A,B,andPA,B,\ $\text${and}\ P i.e., $D\in\mathrm{conv}\left\{A,B,P\right\}$.
@@
Figure 4. An illustration of Lemma 3.13: The textured regions (honeycomb and hatched) denote the respective possible locations for $D$ and $P$. Correspondingly, $P$ is contained in the region(dotted) where it satisfies $D\in\mathrm{conv}\left\{A,B,P\right\}$.
Since the segments $\overline{AB}$ and $\overline{CD}$ intersect, and no three points are collinear, the points $C$ and $D$ lie on either side of the line $L_{AB}$; assume without loss of generality, $C\in L^{-}_{AB}$ and $D\in L^{+}_{AB}$. Further, the point $D$ being closer to $C$ than both $A$ and $B$ implies $D\in L^{-}_{A|C}\cap L^{+}_{C|B}$. Likewise, $P\in L^{+}_{A|C}\cap L^{-}_{C|B}$. Now it can be easily checked that $L^{+}_{A|C}\cap L^{-}_{C|B}\subset L^{+}_{AD}\cap L^{-}_{BD}$ i.e., $P\in L^{+}_{AD}\cap L^{-}_{BD}$ (See Figure 4 for an illustration). Consequently, $D\in\mathrm{conv}\left\{A,B,P\right\}$ as follows from Lemma 3.12.
@@
Let $\delta>0$. Suppose A,B,C,D,X,YA,~B,~C,~D,~X,~Y are six points in $\mathbb{R}^{2}$ satisfying the following:
@@
    ABCDABCD is a quadrilateral with each side of length less than $\delta$ and each diagonal of length at least $\delta$.
@@
    The distance from $X$ to $C$ and $D$ is less than $\delta$, and the distance from $Y$ to $A$ and $B$ is less than $\delta$.
@@
    The distance between $X$ and $Y$ is at least $\delta$.
Then, $X$ is a guard point of $ABCD$ along the side $AB$ if and only if $Y$ is a guard point of $ABCD$ along the side $CD$. Furthermore, in either case, both points $X$ and $Y$ are less than $\delta$ away from $A$, $B$, $C$, and $D$.
@@
Figure 5. An illustration of Lemma 3.14. In this arrangement, the point $Y$ lies in the intersection of hatched and dotted regions. (Right) The dotted region represents all guard points of $ABCD$ along the side CDCD; the segments $\overline{YA}$ and $\overline{YB}$ intersect $\overline{CD}$.
We prove one direction; the converse follows analogously by swapping A,BA,~B with C,DC,~D.
Suppose $X$ is a guard point of $ABCD$ along the side $AB$. From $(1)$, it follows that the quadrilateral $ABCD$ is simple and convex. The points $A$, $B$, and $X$ are not collinear as follows from $(2)$ and $(3)$. Hence, without loss of generality, assume the points $C$ and $D$ are on the left side, and $X$ on the right side of the line $L_{AB}$. The point $Y$ being closer to both $A$ and $B$ than $X$ implies $Y\in L^{+}_{A|X}\cap L^{-}_{X|B}$ (An illustration of a possible arrangement is shown in Figure 5). We have the following sequential observations:
  1. (a)
    $C\in L^{-}_{A|X}$ since the point $C$ is closer to $X$ than $A$. Similarly, $D\in L^{+}_{X|B}$ since the point $D$ is closer to $X$ than $B$.
  2. (b)
@@
  3. (c)
    However, $Y\notin\mathrm{conv}\{A,B,C,D,X\}$ since $d(X,Y)\geq\delta$. Hence, $Y\in L^{-}_{CD}$.
  4. (d)
    Additionally, $\mathrm{(a)}$ also implies $L^{-}_{CD}\cap(L^{+}_{A|X}\cap L^{-}_{X|B})\subset L^{+}_{AC}\cap L^{-}_{BD}$ (See Figure 5(a)).
  5. (e)
    Since $d(B,D)\geq\delta$, it follows that $D\notin\mathrm{conv}\left\{A,B,Y\right\}$. Consequently, by Lemma 3.12, we obtain $Y\in L^{-}_{AD}\cup L^{+}_{BD}$. Similarly, we have $Y\in L^{-}_{AC}\cup L^{+}_{BC}$.
Henceforth, from the observations $(\mathrm{d})$ and $(\mathrm{e})$, it follows that $Y\in L^{-}_{CD}\cap L^{+}_{BC}\cap L^{-}_{AD}$, the region of all guard points of $ABCD$ along the side $CD$ (See Figure 5(b)). More precisely, $Y$ is a guard point along the side $CD$.
Furthermore, since the segments $\overline{YA}$ and $\overline{YB}$ intersects the segment $\overline{CD}$, it follows from Lemma 2.4 that the distance of $Y$ from points $C$ and $D$ is less than $\delta$. Likewise, the distance of $X$ from points $A$ and $B$ is less than $\delta$.
Note that the planar-Rips complex induced by the six points in the above lemma is nothing but the boundary complex of octahedron $\mathcal{O}_{2}$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]